OpenStudy (impracticaljoker):

Is anyone good at 8th grade math? I'd really like some help.

jabez177 (jabez177):

Post the question. :P

OpenStudy (impracticaljoker):

A student is trying to solve the system of two equations given below: Equation P: a + b = 6 Equation Q: 4a + 2b = 19 Which of the following steps can be used to eliminate the a term? (4 points) –1(4a + 2b = 19) –4(4a + 2b = 19) –4(a + b = 6) 4(a + b = 6)

OpenStudy (impracticaljoker):

It's A right?

jabez177 (jabez177):

Yeah.

OpenStudy (impracticaljoker):

Jordan wants to solve the following system using the elimination method: 2x + 3y = 10 x + y = 7 What number should the equation x + y = 7 be multiplied by to eliminate y? (4 points) –2 2 –3 3

OpenStudy (impracticaljoker):

@jabez177